OpenStudy (bezerkhd):

Below are the steps to solve an equation: Step 1: |x − 5| + 2 = 5 Step 2: |x − 5| = 5 − 2 Step 3: |x − 5| = 3

OpenStudy (bezerkhd):

Possible Answers: x + 5 = −3 −x − 5 = 3 x − 5 = −3

OpenStudy (bezerkhd):

I've already rules out D

OpenStudy (bezerkhd):

@Preetha could you help?

pooja195 (pooja195):

Step 1: |x − 5| + 2 = 5 Step 2: |x − 5| = 5 − 2 Step 3: |x − 5| = 3 Answer choices: x + 5 = −3 −x − 5 = 3 x − 5 = −3 correct?

OpenStudy (bezerkhd):

yep

pooja195 (pooja195):

What is the question?

OpenStudy (bezerkhd):

Which of the following is a correct next step to solve the equation?

pooja195 (pooja195):

Ok so are aware of the fact that an absolute value can have a negative or positive SOLUTION?

OpenStudy (bezerkhd):

yep

pooja195 (pooja195):

Knowing that what would our answer be? Keep in mind you simply remove the absolute value bars you don't change anything there.

OpenStudy (bezerkhd):

I've gotten this question wrong 2 times already and i'm trying to get it right this time

OpenStudy (bezerkhd):

k

pooja195 (pooja195):

If you feel that i'm wrong (when I haven't even heard an answer from you) then I think someone else can help

OpenStudy (bezerkhd):

No, i'm not saying you're wrong.

OpenStudy (bezerkhd):

I'm saying that I've been making the same mistake and I want to know what it is

pooja195 (pooja195):

Let's move on from this and just finish up the problem. Knowing that what would our answer be? Keep in mind you simply remove the absolute value bars you don't change anything there.

OpenStudy (bezerkhd):

WOW.... face palm*

OpenStudy (bezerkhd):

well sorry about that.

OpenStudy (bezerkhd):

im dumb -_-

pooja195 (pooja195):

I'd like to see a bit more effort and confidence out of you before I help.

OpenStudy (bezerkhd):

C right?

pooja195 (pooja195):

Nope Keep in mind you simply remove the absolute value bars you don't change anything there. "-x" is not supposed to be -x

pooja195 (pooja195):

Try again

OpenStudy (bezerkhd):

Yea i'm stumped.

pooja195 (pooja195):

Ok can you just show steps as to how you would solve this? Step 1: |x − 5| + 2 = 5

OpenStudy (bezerkhd):

x-5+2=5 -2 -2 x-5=5 +5 +5 x=10

pooja195 (pooja195):

Ok good now remember how I said that absolute values can also have a negative solution? What do we do with the 5 on the right side?

OpenStudy (bezerkhd):

Assuming you mean the original equation, make the solution negative .

pooja195 (pooja195):

Perfect :D

pooja195 (pooja195):

So our answer would be...?

OpenStudy (bezerkhd):

Think its 'A' though i'm not 100%

pooja195 (pooja195):

Look at your equation again 

OpenStudy (bezerkhd):

B

pooja195 (pooja195):

Wait i'm sorry >.< C was right

OpenStudy (bezerkhd):

lol

pooja195 (pooja195):

I didn't see that you said that .-.

OpenStudy (bezerkhd):

b was the only one left so i said b

OpenStudy (bezerkhd):

its fine

pooja195 (pooja195):

I know but C is correct you had the right idea then, do you understand?

OpenStudy (bezerkhd):

I just wanted to know how to do it so mission accomplished!

OpenStudy (bezerkhd):

and yea thanks!

pooja195 (pooja195):

yay! :) you're welcome ^_^